### v1

# references.bib

Abstract
A theorem by Ding, Oporowski, Oxley, and Vertigan states that every sufficiently large bipartite graph without twins contains a matching, co-matching, or half-graph of any given size as an induced subgraph. We prove that this Ramsey statement has polynomial dependency assuming bounded VC-dimension of the initial graph, using the recent verification of the Erdős-Hajnal property for graphs of bounded VC-dimension. Since the theorem of Ding et al. plays a role in (finite) model theory, which studies even more restricted structures, we also comment on further refinements of the theorem within this context.
1 Introduction
Ramsey theory is an area of mathematics asserting that each sufficiently large object contains a still large well-structured subobject. The foundational theorem, proved by Frank Ramsey in 1930, states that every sufficiently large graph contains a still large homogeneous subgraph, i.e. a clique or an independent set [Ramsey1930]. It is well known by a theorem of Erdős and Szekeres [Erdos1935] and a probabilistic construction by Erdős [Erdos1947] that there is always a logarithmically large homogeneous set, but we cannot hope to find a larger one in general. Finding the exact base of the logarithm is one of the most difficult and important problems in Ramsey theory, which have recently witnessed a major breakthrough by the work of Campos, Griffiths, Morris, and Sahasrabudhe [Campos2023], which was subsequently optimized by Gupta, Ndiaye, Norin, and Wei [Gupta2024].
To obtain a bigger than logarithmically large homogeneous set, it is necessary to put additional assumptions on the given graphs. A class of graphs $\mathcal{C}$ is said to have the Erdős-Hajnal property if there is a constant $e>0$ such that each graph $G\in\mathcal{C}$ on $n$ vertices has a homogeneous subset of size $n^{e}$. The celebrated Erdős-Hajnal conjecture states that for each graph $H$ the class of $H$-free graphs admits the Erdős-Hajnal property [Erdos1977]. The conjecture is widely open in full, and was only recently resolved for all graphs on $5$ vertices [Nguyen2023][Chudnovsky2013].
A more restrictive property for a graph class than being $H$-free is having bounded VC-dimension. The Vapnik–Chervonenkis dimension, VC-dimension for short, is a measure of complexity of combinatorial structures, originally invented in the context of learning theory [Vapnik1971], which later proved to be fruitful in combinatorics as well. Structures with bounded VC-dimension often possess better Ramsey properties than general graphs [Fox2021][Janzer2024][Balogh2024]. In particular, Nguyen, Scott, and Seymour recently proved that graphs of bounded VC-dimensions admit the Erdős-Hajnal property [NSS_2023].
In this paper, we consider the following Ramsey statement by Ding, Oporowski, Oxley, and Vertigan [Ding_1996], independently proved using a similar method by Gravier, Maffray, Renault, and Trotignon [Gravier2004] building upon a result of Bauslaugh [Bauslaugh2001].
Theorem 1 ([Ding_1996][Gravier2004]).
There is a monotone unbounded function $f:\mathbb{N}\to\mathbb{N}$ such that the following holds. Let $G=(U,V,E)$ be a bipartite graph such that the part $B$ has $n$ vertices and no twins. Then it contains an induced subgraph $F=(U^{\prime},V^{\prime},E^{\prime})$ with $|U^{\prime}|=|V^{\prime}|\geq f(n)$ of $G$ that is isomorhic to either:
  1. (i)
    a matching,
  2. (ii)
    a co-matching,
  3. (iii)
    a half-graph.
Here the term twins refers to a pair of vertices with the same set of neighbors. All the graphs from the conclusion are bipartite with parts $u_{1},\dots,u_{n}$ and $v_{1},\dots,v_{n}$. The matching is the graph with $u_{i}\sim v_{j}$ iff $i=j$. The co-matching is the bipartite complement of a matching, i.e. $u_{i}\sim v_{j}$ iff $i\not=j$. The half-graph is the graph with $u_{i}\sim v_{j}$ iff $i\leq j$. Note that the binary complement, i.e. $u_{i}\sim v_{j}$ iff $i>j$, of the half-graph of order $n$ contains a half-graph of order $n-1$ as an induced subgraph.
Ding et al. proved Theorem 1 with log-log dependency of the size of $H$ on the size of $G$, and Gravier et al. showed that we cannot hope for better than logarithmic dependency. Inspired by the recent progress on the Erdős-Hajnal property, Dreier and Toruńczyk asked, [^1] In an open problem session within Algomanet Workshop in September 2024, Warsaw. whether Theorem 1 holds with a polynomial bound under the additional assumption that the graph $G$ has bounded VC-dimension.
Here we confirm that this is indeed true. Our main result is the following theorem.
Theorem 2.
For every $d\in\mathbb{N}$, there is $c>0$ such that the following holds. Let $G=(U,V,E)$ be a bipartite graph of VC-dimension at most $d$ such that the part $V$ has $n$ vertices and no twins. Then it contains an induced subgraph $F=(U^{\prime},V^{\prime},E^{\prime})$ with $|U^{\prime}|=|V^{\prime}|\in\Omega_{d}(n^{c})$ that is isomorhic to either:
  1. (i)
    a matching,
  2. (ii)
    a co-matching,
  3. (iii)
    a half-graph.
Moreover, since Theorem 1 plays a role in model theory, where it is often considered to be a folklore result [Bonnet2024][Gajarsky2023], we comment on further improvements of Theorem 2 within model-theoretic context in Section 4. In fact, the applications in model theory was the original motivation of this paper.
2 Preliminaries
All our graphs are simple and finite. All our matrices are binary. Following the paper of Ding et al. [Ding_1996], we mostly work in the language of biadjacency matrices. The biadjacency matrix of a bipartite graph $G=(U,V,E)$ is the matrix $A\in\{0,1\}^{|U|\times|V|}$ with $A_{i,j}=1$ iff $\{u_{i},v_{j}\}\in E$. Natural operations for biadjacency matrices include permutation and deletion of rows/columns. Note that a matrix $B$ that is obtained from $A$ by these operations is the biadjacency matrix of an induced subgraph of $G$. We call such a matrix $B$ a submatrix of $A$.
Given values $\alpha,\beta,\gamma\in\{0,1\}$, we say that a matrix $A\in\{0,1\}^{m\times n}$ is a $(\alpha,\beta,\gamma)$-matrix if it is of the form:
$A_{i,j}=\begin{cases}\alpha&\text{ if }i<j,\\ \beta&\text{ if }i=j,\\ \gamma&\text{ if }i>j.\end{cases}$
Observe that the square $(\alpha,\beta,\gamma)$-matrices with $\alpha\not=\beta$ correspond to the biadjacency matrices of the graphs from the conclusion of Theorem 2 (and the binary complement of a half-graph). Consequently, our result can be stated in the matrix language as follows.
Theorem 3.
For every $d\in\mathbb{N}$, there is $c>0$ such that the following holds. Let $A$ be a binary matrix of VC-dimension at most $d$ with at least $n$ columns, no two of which are identical. Then it contains a square $(\alpha,\beta,\gamma)$-submatrix with $\alpha\not=\beta$ of size $n^{c}\times n^{c}$.
The VC-dimension of a matrix is defined in the next section.
2.1 VC-dimension
The Vapnik–Chervonenkis dimension is a numeric parameter capturing the complexity of a set system [Vapnik1971].
Definition 1 (VC-dimension of a set system).
Consider a set system $\mathcal{S}=(U,\mathcal{F})$, where $U$ is a finite set and $\mathcal{F}$ is a family of subsets of $U$. We say that $X\subseteq U$ is shattered if
|$\big{|}\{X\cap F:F\in\mathcal{F}\}\big{|}=2^{|X|}.$
The VC-dimension of the set system $\mathcal{S}$, denoted by $\operatorname{VC-dim}(\mathcal{S})$, is the cardinality of the largest shattered subset of $U$.
There are several natural set systems, whose VC-dimension is of our interest. For a graph $G$, we have the system $(V,\{N(v):v\in V\})$ of neighborhoods of $G$. More generally, for a matrix $A\in\{0,1\}^{m\times n}$, we consider the system of columns $([m],\{A_{*,j}:j\in[n]\})$, where the column $A_{*,j}$ stands for the set $\{i\in[m]:A_{i,j}=1\}$. Similarly, the matrix $A$ has an associated system of rows $([n],\{A_{i,*}:i\in[m]\})$. Note that if $A$ is the adjacency matrix of a graph $G$, then the column and row systems of $A$ are the same (as $A$ is symmetric), and they agree with the system of neighborhoods of $G$.
We use these systems to define the VC-dimension of graphs and matrices.
Definition 2 (VC-dimension of a graph).
The VC-dimension of a graph $G=(V,E)$, $\operatorname{VC-dim}(G)$, is defined as the VC-dimension of the system $(V,\{N(v):v\in V\})$ of neighborhoods of $G$.
Definition 3 (VC-dimension of a matrix).
The VC-dimension of a matrix $A\in\{0,1\}^{m\times n}$, $\operatorname{VC-dim}(A)$, is defined as the maximum of the VC-dimensions the system of its columns $([m],\{A_{*,j}:j\in[n]\})$ and the system of its rows $([n],\{A_{i,*}:i\in[m]\})$.
As hinted, the definition of the VC-dimension is consistent for a graph and its adjacency matrix. Furthermore, the definitions also agree for a bipartite graph and its biadjacency matrix.
Observation 1.
Let $G$ be a graph and $A$ its adjacency matrix, then $\operatorname{VC-dim}(G)=\operatorname{VC-dim}(A)$. Moreover, if $G$ is bipartite and $B$ is its biadjacency matrix, then $\operatorname{VC-dim}(G)=\operatorname{VC-dim}(B)$.
VC-dimension is monotone with respect to taking submatrices (i.e. induced subgraphs) and their complements.
Observation 2.
If $B$ is (the binary complement of) a submatrix of $A$ (or its transposition), then $\operatorname{VC-dim}(B)\leq\operatorname{VC-dim}(A)$.
2.2 Growth function
A closely related notion to VC-density is the growth function of a system.
Definition 4 (Growth function of a set system).
Let $\mathcal{S}=(U,\mathcal{F}),|U|=N,$ be a set system. The growth function $\pi_{\mathcal{S}}:[N]\to\mathbb{N}$ is defined as
$\pi_{\mathcal{S}}(n)=\max_{X\subseteq U,|X|=n}\big{|}\{X\cap F:F\in\mathcal{F}\}\big{|}.$
The classical Sauer-Shelah lemma states that systems of bounded VC-dimension have the growth function bounded by a polynomial [Sauer_1972][Shelah_1972].
Theorem 4 (Sauer-Shelah lemma [Sauer_1972][Shelah_1972]).
Let $\mathcal{S}=(U,\mathcal{F})$ be a set system of VC-dimension at most $d$. Then for any set $X\subseteq U$ of size $n$, we have
|$\big{|}\{X\cap F:F\in\mathcal{F}\}\big{|}\leq\Phi_{d}(n)\coloneqq\sum_{i=0}^{d}\binom{n}{i}.$
2.3 Erdős-Hajnal property
As stated in the introduction, we say that a class of graphs $\mathcal{C}$ has the Erdős-Hajnal property if the graphs from $\mathcal{C}$ admit polynomially large homogeneous sets. A recent theorem of Nguyen, Scott, and Seymour that verifies the Erdős-Hajnal property for graphs of bounded VC-dimension [NSS_2023] plays a crucial part in our proof.
Theorem 5 ([NSS_2023]).
For every $d\geq 1$, there exists $e(d)>0$ such that every graph $G$ on $n$ vertices of VC-dimension at most $d$ contains a homogeneous set of size at least $n^{e(d)}$.
3 Proof
Given a matrix of VC-dimension at most $d$ with at least $n$ columns, no two of which are identical, our goal is to find a square $(\alpha,\beta,\gamma)$-submatrix with $\alpha\not=\beta$ of polynomial size. The original proof of Ding et al. is carried out in two steps. First, they obtain a square $(\alpha,\beta,*)$-submatrix of the given matrix with $\alpha\not=\beta$, where * is a Joker (i.e. no restriction for entries below the diagonal). As the second step, they apply the Ramsey theorem to an auxiliary graph to homogenize the lower triangular matrix. Both steps shrink the matrix logarithmically.
In our case, we prove that the auxiliary graph constructed from a matrix of bounded VC-dimension has bounded VC-dimension as well (Lemma 2). This allows us to perform the second step from the $(\alpha,\beta,*)$-submatrix to the $(\alpha,\beta,\gamma)$-submatrix with polynomial dependency by Theorem 5.
Moreover, we show that after a convenient preparatory step, it is possible to use the Ramsey theorem, i.e. in our case Theorem 5, in the first step to obtain the $(\alpha,\beta,*)$-submatrix. Indeed, we initially produce a switch submatrix (Lemma 1), which allows us to choose the value $\beta$ on the diagonal after we obtain the value $\alpha$ above the diagonal from the Ramsey theorem (Lemma 3). Then it remains to homogenize the values below the diagonal as mentioned above (Lemma 4).
3.1 Switch submatrix
We begin by defining the notion of a switch matrix. Then we show that under the assumptions of Theorem 3, the given matrix contains a polynomially large switch submatrix.
Definition 5.
A matrix $A\in\{0,1\}^{n\times 2n}$ is a switch matrix of size $n$ if it is of the following form:
  1. (i)
    for all $i\in[n]$, we have $A_{i,2i-1}=0$ and $A_{i,2i}=1$,
  2. (ii)
    for all $i,j\in[n],i<j$, we have $A_{i,2j-1}=A_{i,2j}$.
The following important technical lemma says that in our setting of bounded VC-dimensions we have polynomially large switch submatrices. We present the lemma in greater generality for matrices whose column systems have the growth function bounded by a polynomial.
Lemma 1.
Let $A$ be a matrix whose system of columns has the growth function $\pi(n)\leq bn^{a}$ with $a\geq 1,b>0$. Suppose $A$ has $n$ columns, no two of which are identical. Then $A$ contains a switch submatrix $B$ of size at least $zn^{1/a}$ for any $z<1/(2+b)$.
Proof.
Starting from the empty submatrix $B_{0}$, we greedily construct switch submatrices $B_{1},B_{2},\dots$, with $B_{i}$ having size $i$. Each $B_{i+1}$ is obtained by extending $B_{i}$ by a new row and two columns. We only need to prove that we can run the process for at least $zn^{1/a}$ steps.
Suppose we have a switch submatrix $B_{i}$ of $A$ of size $i$ that uses rows $\mathcal{R}$ and columns $\mathcal{C}$ from the original matrix (likely in different order). For $x\in\{0,1\}^{\mathcal{R}}$, let $\Gamma(x)$ be the set of columns of $A$ that have the pattern $x$ at rows $\mathcal{R}$. To extend the submatrix $B_{i}$, we want to find a vector $x^{*}\in\{0,1\}^{\mathcal{R}}$ such that $|\Gamma(x^{*})\setminus\mathcal{C}|\geq 2$. This is indeed sufficient: take distinct columns $C,C^{\prime}\in\Gamma(x^{*})\setminus\mathcal{C}$. By assumption, these columns differ at some row $R$, which certainly lies outside of $\mathcal{R}$. Thus, we can extend $B_{i}$ by the row $R$ and columns $C,C^{\prime}$ to form the submatrix $B_{i+1}$. To be precise, the last column is the one of $C$ or $C^{\prime}$ that has the value $1$ in the row $R$. Clearly, the resulting submatrix $B_{i+1}$ is then a switch matrix of size $i+1$.
It remains to show that such a vector $x^{*}\in\{0,1\}^{\mathcal{R}}$ exists when extending the submatrix $B_{i}$ with $i<zn^{1/a}$. By the assumption on the growth function, we have at most $bi^{a}$ vectors $x\in\{0,1\}^{\mathcal{R}}$ for which the set $\Gamma(x)$ is non-empty. Therefore, the average size of a non-empty $\Gamma(x)\setminus\mathcal{C}$ is at least
$\frac{n-|\mathcal{C}|}{bi^{a}}=\frac{n-2i}{bi^{a}}>\frac{n-2zn^{1/a}}{b(zn^{1/a})^{a}}=\frac{1}{bz^{a}}\left(1-\frac{2z}{n^{1-1/a}}\right)\geq\frac{1}{bz}(1-2z)>1$
as we need. Note that the last inequality is equivalent with the assumption $z<1/(2+b)$. ∎
We get an immediate corollary for matrices of bounded VC-dimension.
Corollary 1.
Let $A$ be a matrix of VC-dimension at most $d$ with $n$ columns, no two of which are identical. Then $A$ contains a switch submatrix $B$ of size at least $\Omega_{d}(n^{1/d})$.
Proof.
By Theorem 4, the class of matrices of VC-dimension at most $d$ has the growth function bounded by $\Phi_{d}(n)\in\mathcal{O}_{d}(n^{d})$. The conclusion then follows from Lemma 1. ∎
3.2 Upper graph and its VC-dimension
Here we define the notion of an auxiliary graph called upper graph associated to a square matrix $A$. We prove that if $A$ has bounded VC-dimension, the corresponding upper graph has bounded VC-dimension as well.
Definition 6.
Let $A$ be a square matrix of size $n$. We call the graph $G=([n],\{\{i,j\}:i<j,A_{i,j}=1\})$ the upper graph of the matrix $A$.
In other words, the upper graph of the matrix $A$ is the graph $G$ whose adjacency matrix agrees with $A$ in the entries above the diagonal.
Lemma 2.
Let $A\in\{0,1\}^{n\times n}$ be a square $(*,0,*)$-matrix and $G$ be the upper graph of $A$. For any $d\in\mathbb{N}$, we have that $\operatorname{VC-dim}(G)\geq 2d$ implies $\operatorname{VC-dim}(A)\geq d/2$.
Conversely, for any $d\in\mathbb{N}$, if $\operatorname{VC-dim}(A)\leq d$, then $\operatorname{VC-dim}(G)\leq 4d+1$.
Proof.
Let $B$ be the adjacency matrix of $G$. The VC-dimension of $G$ is at least $2d$, which is witnessed by a shattered set $X=\{v_{1}<\dots<v_{2d}\}\subseteq[n]$ of size $d$. Let $Y=\{u_{1}<\dots<u_{2^{2d}}\}\subseteq[n]$ be the set of vertices whose neighborhoods witness that $X$ is shattered.
We divide the set $X$ into $X_{1}=\{v_{1}<\dots<v_{d}\}$ and $X_{2}=\{v_{d+1}<\dots<v_{k}\}$. Moreover, we split the set $Y$ into $Y_{1}=\{u_{i}\in Y:u_{i}<v_{d}\}$ and $Y_{2}=\{u_{i}\in Y:u_{i}\geq v_{d}\}$. Certainly, one of the sets $Y_{1}$ or $Y_{2}$ has size at least $2^{2d-1}$. If $|Y_{2}|\geq 2^{2d-1}$, the submatrix $C$ of $B$ induced by the rows $X_{1}$ and columns $Y_{2}$ lies entirely in the upper triangle of $B$, so it is a submatrix of $A$ as well. In the other case, if $|Y_{1}|\geq 2^{2d-1}$, the submatrix $D$ of $B$ induced by $X_{2}\times Y_{1}$ lies in the bottom triangle of $B$. By symmetry of $B$, the transposition of $D$ appears in the upper triangle of $B$ and hence also in $A$. The rest of the proof is then the same with the submatrix $D^{\top}$ instead of $C$ and with the role of rows and columns reversed.
[…]
$2^{d-1}\leq\text{ \#distinct patterns on $X_{1}$ }\leq\Phi_{\operatorname{VC-dim}(A)}(d)=\sum_{i=0}^{\operatorname{VC-dim}(A)}\binom{d}{i},$
which implies that $\operatorname{VC-dim}(A)\geq d/2$.
The part “Conversely” is the contraposition with correct rounding that includes odd values of $\operatorname{VC-dim}(G)$d-1. ∎
3.3 Homogenization
In this part, we utilize Theorem 5 to first extract an $(\alpha,\beta,*)$-matrix with $\alpha\not=\beta$ from a switch matrix and then an $(\alpha,\beta,\gamma)$-matrix from the $(\alpha,\beta,*)$-matrix. The polynomial dependency in Theorem 5 ensures polynomial dependency in our statements.
Lemma 3.
Let $A$ be a switch matrix of size $n$ of VC-dimension at most $d$. Then it contains a square $(\alpha,\beta,*)$-matrix with $\alpha\not=\beta$ of size at least $n^{e(4d+1)}$, where $e(\cdot)$ is the function from Theorem 5.
Proof.
First, we consider the square $(*,0,*)$-submatrix $B$ of $A$ obtained by deleting all even columns of $A$. Let $G$ be the upper graph of $B$. By Observation 2 and Lemma 2, the VC-dimension of $G$ is at most $14d+$ 1. Theorem 5 then implies that $G$ contains a homogeneous subset of vertices $X$ of size at least $n^{e(4d+1)}$. We consider the submatrix $C$ of $A$ with rows $\mathcal{R}=X$, and columns $\mathcal{C}=\{2i-1:i\in X\}$ if $G[X]$ is a clique, resp. $\mathcal{C}=\{2i:i\in X\}$ if $G[X]$ is an independent set. Clearly, $C$ is a square matrix of size at least $n^{e(4d+1)}$.
We claim that the matrix $C$ is an $(\alpha,\beta,*)$-matrix with $\alpha\not=\beta$. If $G[X]$ is a clique, then for each $i,j\in X,i<j$, we have that $A_{i,2j-1}=1$. Moreover, by definition of a switch matrix, we have for all $i\in X$ that $A_{i,2i-1}=0$. Therefore, the matrix $C$ is a $(1,0,*)$-matrix.
If $G[X]$ is an independent set, then for each $i,j\in X,i<j$, we have that $A_{i,2j-1}=0$. Thus, also $A_{i,2j}=0$ by definition of a switch matrix. Moreover, we have for all $i\in X$ that $A_{i,2i}=1$. Therefore, the matrix $C$ is a $(0,1,*)$-matrix. ∎
Lemma 4.
Let $A$ be a square $(\alpha,\beta,*)$-matrix of size $n$ of VC-dimension at most $d$. Then it contains a square $(\alpha,\beta,\gamma)$-matrix with $\alpha\not=\beta$ of size at least $n^{e(4d+1)}$, where $e(\cdot)$ is the function from Theorem 5.
Proof.
We proceed similarly as in the proof of Lemma 3, except that we construct the upper graph from the transposition of the matrix $A$, resp. from the the binary complement of the transposition if $A$ is a $(0,1,*)$-matrix (to formally complain with Lemma 2 by having 00 on the diagonal). By applying Theorem 5, we obtain a square $(\alpha,\beta,\gamma)$-submatrix of $A$ with $\alpha\not=\beta$ of size at least $n^{e(4d+1)}$. ∎
It remains to prove the main theorem.
Proof of Theorem 3.
By composing Corollary 1 with Lemmas 3 and 4, we obtain an $(\alpha,\beta,\gamma)$-diagonal submatrix with $\alpha\not=\beta$ of polynomial size with the exponent
$\frac{1}{d}\cdot e(4d+1)\cdot e(4d+1),$
where $e(\cdot)$ is the function from Theorem 5. ∎
4 Remarks on tame classes
In recent years, much effort has been invested into combinatorial understanding of classical model-theoretic notions such as stability and dependence with the aim to establish dividing lines between tame and wild classes of finite structures [Gajarsky2023][Torunczyk2023][Dreier2024]. These include the model-theoretic notions of (monadic) stability, (monadic) dependence and combinatorial notion of bounded twin-width. For their definitions please see [Tent2012][Simon2015][Dreier2024][Bonnet2021], Figure 1 shows their inclusions. Considered folklore, Theorem 1 plays an important role in this area. For example, it was one of the tools that allowed Dreier, Mählmann, and Toruńczyk to characterize the forbidden subgraphs of monadically dependent graph classes, yielding the first purely combinatorial definition of monadic dependence [Dreier2024].
Figure 1: A diagram of various tameness notions and their inclusions.
The coarsest of these notions, dependence, implies bounded VC-dimension. Therefore, our improvement in form of Theorem 2 applies for dependent classes of graphs and for all classes from Figure 1. However, the more restricted classes allow for further improvements of Theorem 2 and better bounds.
In Corollary 1, we used Theorem 4 to bound the growth function of classes of bounded VC-dimension. While Theorem 4 gives a polynomial bound on the growth function of the neighborhood systems, also called neighborhood complexity, it is conjectured that the monadically dependent classes admit an almost linear bound [Dreier2023a][Dreier2024]. That is, that for graphs from the class $\mathcal{C}$, the growth function is bounded by $bn^{1+\varepsilon}$ for every $\varepsilon>0$ so some constant $b=b(\varepsilon,\mathcal{C})$. This is known to be true for monadically stable classes [Dreier2023b]. For classes of bounded twin-width, the situation is even more pleasant as the neighborhood complexity is linear as proved independently by [Bonnet2022] and [Przybyszewski2023], with an essentially tight bound by [Bonnet2024a]. Using these exponents in Lemma 1, we conclude that graphs from classes of bounded twin-width admit a switch submatrix of a linear size, while monadically stable classes admit a polynomial switch submatrix with the exponent $1-\varepsilon$ for every $\varepsilon>0$.
Moreover, if the upper graph of the given matrix belongs to a tame class, we can reach better bounds by substituting Theorem 5 by a more specific statement. It was implicitly proved by Malliaris and Shelah [Malliaris2013, Theorem 3.5] that stable graph classes posses the Erdős-Hajnal property. The bound was further improved in [Malliaris2017, Corollary 4.9] using only the milder assumption that the graph class is edge-stable (that is, only the edge relation is required to be stable; not necessarily all the definable relations). Furthermore, for the more restricted monadically stable graph classes, it was recently proved that they possess the Erdős-Hajnal property with the exponent $1/2-\varepsilon$ for any $\varepsilon>0$ [Braunfeld2025, Corollary 62]. Notably, this says that there is a uniform Erdős-Hajnal exponent for all monadically stable graph classes.
Speaking about stable classes, we should point out that we can further simplify the conclusion of Theorem 2. Indeed, for each stable graph class there is a bound on the size of half-graphs that can appear as subgraphs, which allows us to narrow the conclusion of Theorem 2 that the target subgraph is either a matching or a co-matching.
5 Acknowledgement
The author would like to thank Patrice Ossona de Mendez for his many suggestions concerning Section 4.
The author is partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC Synergy Grant DYNASNET, grant agreement No 810115), and by Project 24-12591M of the Czech Science Foundation (GAČR).
[…]